Your firm has a cost of goods sold of $75 thousand dollars, an average inventory of $26 thousand dollars, credit sales of $36 thousand, and average recievables of $13 thousand. What is the length of your firm's operating cycle?
Answer #1
Inventory period = ( Average Inventory / COGS ) * Days in a year = ( 26000 / 75000 ) * 365 127 Days
Accounts receivables period = ( Average Accounts receivables / Credit Sales ) * Days in a year = ( 13000 / 36000 ) * 365 132 days
Operating cycle = Inventory period + Accounts receivables period = 127 + 132 259 Days